OpenStudy (anonymous):

What are the solutions for x in the proportion x + 6 over 2x = x over 2

OpenStudy (anonymous):

\[\frac{x+6}{2x}=\frac{x}{2}?\] Cross multiply: \[2(x+6)=x(2x)\] simplify and solve for x. Let us know how far you get.